What affect the brightness of two identical bulbs .and will they be brighter in series or parallel series
1
Expert's answer
2019-05-17T11:17:47-0400

The brightness of the lamps is proportional to the power dissipation

With parallel connection power dissipation is:



"P_{par}=U^2*(R_1+R_2)\/(R_1*R_2)"

With series connection power dissipation is:


"P_{ser}=U^2\/(R_1+R_2)"

Since the lamp resistances are equal, then:


"R_1=R_2"

The difference in power and brightness is:


"P_{par}\/P_{ser}=(2R)^2\/R^2=4"

The brightness of the bulbs is affected by the current passing through the nichrome thread. If the bulbs with the same voltage are connected in series, then the total brightness of the bulbs will decrease by 4 times